# Wasserstein Convergence of Score-based Generative Models under Semiconvexity and Discontinuous Gradients

### v1

@@ 3.3 Main Results - Optimal Data Dimensional Dependence and Rate of Convergence
$\begin{split}W_{2}(\mathscr{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}})\leq C_{1}\sqrt{\epsilon}+C_{2}e^{-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t-\epsilon}+C_{3}(T,\epsilon)\sqrt{\varepsilon_{\text{SN}}}+C_{4}(T,\epsilon)\gamma^{1/2},\end{split}$ (24)
where $C_{1}$, $C_{2}$, $C_{3}$ and $C_{4}$ are given explicitly in Table 2 (Appendix E), $\beta^{\text{OS},K,\mu}_{t}$ is defined in (22), and its integral is computed in Proposition 3.10. In addition, the result in (24) implies that for any $\delta>0$, if we choose $0<\epsilon<\epsilon_{\delta}$, $T>T_{\delta}$, $0<\varepsilon_{\text{SN}}<\varepsilon_{\text{SN},\delta}$ and $0<\gamma<\gamma_{\delta}$ with $\epsilon_{\delta}$, $T_{\delta}$, $\varepsilon_{\text{SN},\delta}$, and $\gamma_{\delta}$ given in Table 2, then
$W_{2}(\mathscr{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}})<\leavevmode\nobreak\ \delta.$
Remark 3.13.
The constant $C_{4}(T,\epsilon)$ in the error bound in (24) contains the optimal dependence of the data dimension, i.e. $O(\sqrt{d})$, which has been found under the more strict assumption of strong-log concavity of $\pi_{\mathsf{D}}$ in [7, Theorem 1 and Remark 12]. However, the optimal dependence of the dimension is achieved at the expenses of a worst rate of covergence of order $1/2$.
The optimal rate of convergence of order $\alpha\in[\frac{1}{2},1]$ for the Euler or Milstein scheme of SDEs with constant diffusion coefficients can be achieved in Theorem 3.12 if we require that $\mathbb{E}[|\hat{\theta}|^{4}]<\infty$ in Assumption 1 and replace Assumption 3.a with Assumption 3.b as stated in Theorem 3.14 below.
Theorem 3.14.
@@
$\begin{split}W_{2}(\mathscr{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}})\leq C_{1}\sqrt{\epsilon}+C_{2}e^{-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t-\epsilon}+C_{3}(T,\epsilon)\sqrt{\varepsilon_{\text{SN}}}+\widetilde{C}_{4}(T,\epsilon)\gamma^{\alpha},\end{split}$ (25)
where $C_{1}$, $C_{2}$, $C_{3}$ and $\widetilde{C}_{4}$ are given explicitly in Table 2 (Appendix E), $\beta^{\text{OS},K,\mu}_{t}$ is defined in (22), and its integral is computed in Proposition 3.10. In addition, the result in (25) implies that for any $\delta>0$, if we choose $0<\epsilon<\epsilon_{\delta}$, $T>T_{\delta}$, $0<\varepsilon_{\text{SN}}<\varepsilon_{\text{SN},\delta}$ and $0<\gamma<\widetilde{\gamma}_{\delta}$ with $\epsilon_{\delta}$, $T_{\delta}$, $\varepsilon_{\text{SN},\delta}$, and $\widetilde{\gamma}_{\delta}$ given in Table 2, then
$W_{2}(\mathscr{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}})<\leavevmode\nobreak\ \delta.$ (26)
Remark 3.15.
In the explicit expression of $\widetilde{C}_{4}$ in Table 2 (Appendix E), the dependence of the data dimension is $O(d)$ due to numerical techniques from [7, Proof of Theorem 10] used in the proof of Theorem 3.14 to achieve the optimal rate of convergence of order $\alpha\in[\frac{1}{2},1]$.
@@ 3.4 Examples of potentials satisfying by Assumption 2
We present several examples to demonstrate the wide applicability of our Assumption 2 to a broad class of data distributions, some of which are not covered by previous results in Wasserstein distance of order two [21, 7, 20, 58, 60, 71].
@@ 3.4.1 Symmetric modified half-normal distribution
where $\Gamma(n)$ is the Gamma function. We refer the reader to Appendix D for additional details about the derivation of (27). As highlighted in [59, Section 2], the modified half-normal distribution appears in several Bayesian statistical methods as a posterior distribution to sample from in Bayesian Binary regression, analysis of directional data, and Bayesian graphical models.
@@
which shows that Assumption 2-(ii) is verified for any $K\geq 0$, and Assumption 2-(iii) is verified for $\mu=2\xi$. Therefore, we can conclude that (27) satisfies Assumption 2.
@@ 3.4.2 Multidimensional Gaussian mixture distribution
We consider a multidimensional Gaussian mixture data distribution with unknown mean and variance, i.e.
$\pi_{\mathsf{D}}(\mathrm{d}x)=\sum_{j=1}^{J}\widetilde{\xi}_{j}\frac{1}{(2\pi\sigma_{j}^{2})^{d/2}}\exp\left(-\frac{|x-\eta_{j}|^{2}}{2\sigma_{j}^{2}}\right)\ \mathrm{d}x,\quad x\in\mathbb{R}^{d},$ (29)
with $\sigma_{j}>0$, $\eta_{j}\in\mathbb{R}^{d}$, and $\widetilde{\xi}_{j}\in[0,1]$ for $j\in\{1,\dots,J\}$ such that $\sum_{j=1}^{J}\widetilde{\xi}_{j}=1$. The authors in [21, Appendix A] show that the score function of (29) is Lipschitz continuous and $-\log\pi_{\mathsf{D}}$ is weakly convex. Therefore, Assumption 2 is satisfied. In addition, the distribution (29) covers also case of the double-well potential:
$U(x)=x^{4}-|x|^{2},\quad x\in\mathbb{R}^{d},$ (30)
@@ 3.4.3 Multi-dimensional Potentials
verifies Assumption 2 with $K=0$, $R=1$, and $\mu=2$ as well as the following non-convex potential presented in [31, Example 4.2]:
@@ 4 Related Work and Comparison
The first approach focuses on $\alpha$-divergences, particularly the Kullback–Leibler (KL) divergence and Total Variation (TV) distance (e.g., [4, 15, 68, 39, 5, 17, 37, 40, 38, 9, 12, 48, 42, 67]), which are the vast majority of the results available in the literature. Crucially, bounds on KL divergence imply bounds on TV distance via Pinsker’s inequality, strengthening their wide applicability. We provide a brief and selective overview of some of the findings following this first approach. The results in TV distance in [37] and in KL divergence [68] assumed that the data distribution satisfies a logarithmic Sobolev inequality and the score function was Lipschitz resulting in convergence bounds characterized by polynomial complexity. By replacing the requirement that the data distribution satisfies a functional inequality with the assumption that $\pi_{\text{D}}$ has finite KL divergence with respect to the standard Gaussian and by assuming that the score function for the forward process is Lipschitz, the authors in [12] managed to derive bounds in TV distance which scale polynomially in all the problem parameters. By requiring only the Lipschitzness of the score at the initial time instead of the whole trajectory, the authors in [9, Theorem 2.5] managed to show, using an exponentially decreasing then linear step size, convergence bounds in KL divergence with quadratic dimensional dependence and logarithmic complexity in the Lipschitz constant. In the work by [4], the authors provide convergence bounds in KL divergence that are linear in the data dimension, up to logarithmic factors, using early stopping and assuming finite second moments of the data distribution. However, both results in [9, Theorem 2.5] and [4, Theorem 1 and Corollary 1] still require the uniqueness of solutions for the backward SDE (3), and therefore additional assumptions on the score function are needed. We refer the reader to [7, Section 4.2] for additional details about this point. Assuming the finiteness of the second moment of the data distribution and using an EI discretization scheme with constant and exponentially decreasing step sizes, the authors in [15, Corollary 2.4] derive a KL divergence bound with early stopping, which scales linearly in the data dimension up to logarithmic factors. Bounds in KL without early stopping have been derived in [15] for data distributions with finite Fisher information with respect to the standard Gaussian distribution. We note that this assumption stated in [15, Assumption H2] still requires the potential $U\in C^{1}(\mathbb{R}^{d})$. The bounds provided in [15, Theorem 2.1 and 2.2] scale linearly in the Fisher information when an EI discretization scheme with constant step size is used and logarithmically in the Fisher information when an exponential-then-constant step size [15, Theorem 2.3] is chosen.
@@
The second approach focuses on convergence bounds in Wasserstein distance, a metric which is often considered more practical and informative for estimation tasks. We can relate results following this approach with the results of the first approach only when $\pi_{\mathsf{D}}$ is a strongly log-concave distribution. In this case, $W_{2}$-bounds in terms of KL divergence follow from an extension of Talagrand’s inequality [23, Corollary 7.2]. However, for two general data distributions, there is no known relationship between their KL divergence and their $W_{2}$. Therefore, we cannot compare our findings in Therem 3.12 and Theorem 3.14 with the results derived following the first approach. One line of work within the second approach assumes (at least) strong log-concavity of the data distribution [20, 7, 60, 58, 71]. Under this (strict) assumption, [7, Remark 12] achieved optimal data dimensional dependence, i.e. reaching $O(\sqrt{d})$. The recent bound in [21, Theorem D.1] scales linearly in the data dimension while relaxing the strong log-concavity assumption on $\pi_{\mathsf{D}}$ to weakly log-concavity, but still requiring that the potential $\nabla^{2}U$ exists (see, e.g., [21, Proof of Proposition B.1 and B.2]). Our Assumption 2 is much weaker than this requirement and it allows to consider the case of potentials with discontinuous gradients covering a wider range of distributions as outlined in Section 3.4. Another line of work following this approach focuses on specific structural assumptions of the data distribution. For instance, convergence bounds in Wasserstein distance of order one with exponential dependence on the problem parameters have been obtained in [6] under the so-called manifold hypothesis, namely assuming that the target distribution is supported on a lower-dimensional manifold or is given by some empirical distribution. Under the same metric, the authors in [46] provide a convergence analysis when the data distribution is defined on a torus. We summarize in Table 1 and the best results obtained in $W_{2}$, i.e. [7, 21] and compare with our best result, which scale polynomially in the data dimension, i.e. $O(\sqrt{d})$ (Theorem 3.12).
@@
Table 1: Summary of previous bounds for $W_{2}(\mathscr{L}(\widehat{Y}_{K}^{\text{EM}}),\pi_{\mathsf{D}})$ and our result in Theorem 3.12. All the bounds assume that $\pi_{\text{D}}(\mathrm{d}x)\propto e^{-U(x)}\mathrm{d}x$ has finite second moments.
@@
$U$ strongly convex, $\nabla\log p_{t}(0)\in L^{2}([\epsilon,T])$, and Assumption 4 $O(\sqrt{d})\sqrt{\epsilon}+O(\sqrt{d})e^{-2\widehat{L}_{\text{MO}}(T-\epsilon)-\epsilon}+O(e^{(1+\zeta-2\widehat{L}_{\text{MO}})(T-\epsilon)})\sqrt{\varepsilon_{\text{SN}}}+O(\sqrt{d}e^{T^{2\alpha+1}}T^{2\alpha+1}\widetilde{\varepsilon}^{1/2}_{\text{AL}})\gamma^{1/2}$, with $\widehat{L}_{\text{MO}}>0$ lower bound of the strongly convex constant of $U$, see e.g., [7, Remark 4]. [7, Remark 12]
$U\in C^{2}(\mathbb{R}^{d})$, weakly convex, and Assumption 4 $e^{(2L_{U}+5)\eta(\beta,L,\gamma)}[e^{-T}W_{2}(\pi_{\mathsf{D}},\pi_{\infty})+4\varepsilon_{\text{SN}}(T-\eta(\beta,L,0))+\sqrt{2\gamma}(4L_{U}d+6d+\sqrt{d+\mathbb{E}[|X_{0}|^{2}]})(T-\eta(\beta,L,0))]$, with $L_{U}\geq 0$ one-sided Lipschitz constant for $\nabla U$, see e.g., [21, Assumption H1], $\eta(\beta,L,\gamma)$ defined in [21, Equation (29)], and $\gamma<2/(2L_{U}+5)^{2}$ [21, Theorem D.1]
Assumption 2 and Assumption 4 $O(\sqrt{d})\sqrt{\epsilon}+O(\sqrt{d})e^{-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t-\epsilon}+O(e^{(1+\zeta)(T-\epsilon)-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t})\sqrt{\varepsilon_{\text{SN}}}+O(\sqrt{d}e^{T^{2\alpha+1}}T^{3\alpha+1}\widetilde{\varepsilon}^{1/2}_{\text{AL}})\gamma^{1/2}$ Theorem 3.12
We close this section by briefly commenting on the choice of deriving our results in Wasserstein distance of order two. Beyond its theoretical relevance, this choice is motivated by practical considerations in generative modeling. First, the Wasserstein distance is often regarded as a more informative and robust metric for estimation tasks. Second, a widely used performance metric for evaluating the quality of images produced by generative models is the Fréchet Inception Distance (FID) [26], which measures the Fréchet distance between the distributions of generated and real samples, assuming Gaussian distributions. In particular, this Fréchet distance is equivalent to the Wasserstein-2 distance. Thus, providing convergence results under the Wasserstein-2 metric enhances the practical relevance of our theoretical findings.

### v2

@@ 3.3 Main Results - Optimal Data Dimensional Dependence and Rate of Convergence
$\begin{split}W_{2}(\mathcal{L}(Y_{J}^{\text{EM}}),\pi_{\mathsf{D}})\leq C_{1}\sqrt{\epsilon}+C_{2}e^{-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t-\epsilon}+C_{3}(T,\epsilon)\sqrt{\varepsilon_{\text{SN}}}+C_{4}(T,\epsilon)\gamma^{1/2},\end{split}$ (26)
where $C_{1}$, $C_{2}$, $C_{3}$ and $C_{4}$ are given explicitly in Table 3 (Appendix E), $\beta^{\text{OS},K,\mu}_{t}$ is defined in 22, and its integral is computed in Proposition 16. In addition, the result in 26 implies that for any $\delta>0$, if we choose $0<\epsilon<\epsilon_{\delta}$, $T>T_{\delta}$, $0<\varepsilon_{\text{SN}}<\varepsilon_{\text{SN},\delta}$ and $0<\gamma<\gamma_{\delta}$ with $\epsilon_{\delta}$, $T_{\delta}$, $\varepsilon_{\text{SN},\delta}$, and $\gamma_{\delta}$ given in Table 3, then
$W_{2}(\mathcal{L}(Y_{J}^{\text{EM}}),\pi_{\mathsf{D}})<~\delta.$
Remark 20.
The constant $C_{4}(T,\epsilon)$ in the error bound in 26 contains the optimal dependence of the data dimension, i.e. $O(\sqrt{d})$, which has been found under the more strict assumption of strong-log concavity of $\pi_{\mathsf{D}}$ in Bruno et al. (2025, Theorem 1 and Remark 12). However, the optimal dependence of the dimension is achieved at the expenses of a worst rate of covergence of order $1/2$.
The optimal rate of convergence of order $\alpha\in[\frac{1}{2},1]$ for the Euler or Milstein scheme of SDEs with constant diffusion coefficients can be attained in Theorem 19 provided that $\mathbb{E}[|\hat{\theta}|^{4}]<\infty$ and that Assumption 3.a is replaced by Assumption 3.b, as stated in Theorem 21 below.
Theorem 21.
@@
$\begin{split}W_{2}(\mathcal{L}(Y_{J}^{\text{EM}}),\pi_{\mathsf{D}})\leq C_{1}\sqrt{\epsilon}+C_{2}e^{-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t-\epsilon}+C_{3}(T,\epsilon)\sqrt{\varepsilon_{\text{SN}}}+\widetilde{C}_{4}(T,\epsilon)\gamma^{\alpha},\end{split}$ (27)
where $C_{1}$, $C_{2}$, $C_{3}$ and $\widetilde{C}_{4}$ are given explicitly in Table 3 (Appendix E), $\beta^{\text{OS},K,\mu}_{t}$ is defined in 22, and its integral is computed in Proposition 16. In addition, the result in 27 implies that for any $\delta>0$, if we choose $0<\epsilon<\epsilon_{\delta}$, $T>T_{\delta}$, $0<\varepsilon_{\text{SN}}<\varepsilon_{\text{SN},\delta}$ and $0<\gamma<\widetilde{\gamma}_{\delta}$ with $\epsilon_{\delta}$, $T_{\delta}$, $\varepsilon_{\text{SN},\delta}$, and $\widetilde{\gamma}_{\delta}$ given in Table 3, then
$W_{2}(\mathcal{L}(Y_{J}^{\text{EM}}),\pi_{\mathsf{D}})<~\delta.$ (28)
Remark 22.
The constant $\widetilde{C}_{4}$, explicitly given in Table 3 (Appendix E) exhibits a linear dependence on the data dimension, i.e., $O(d)$. This scaling arises from the explicit Milstein scheme developed in Kumar & Sabanis (2019), which relies on Assumption 3.b and is leveraged in the proof of Theorem 21 to achieve the optimal convergence rate of order $\alpha\in[\frac{1}{2},1]$. This explicit Milstein scheme requires control on the fourth moment of the one-step discretization, see for instance, Lemma 25 (Appendix C) which enables a convergence rate in $W_{2}$ consistent with the known optimal rate of convergence for the Euler or Milstein scheme of SDEs with constant diffusion coefficients. However, this comes at the cost of a worse dependence on the data dimension.
@@ 3.4 Examples of potentials satisfying by Assumption 2
We present several examples to demonstrate the wide applicability of our Assumption 2 to a broad class of data distributions, some of which are not covered by previous results in Wasserstein distance of order two (Silveri & Ocello, 2025; Strasman et al., 2025; Gao et al., 2025; Bruno et al., 2025; Tang & Zhao, 2024; Yu & Yu, 2025).
@@ 3.4.1 Symmetric modified half-normal distribution
where $\Gamma(n)$ is the Gamma function. We refer the reader to Appendix D for additional details about the derivation of 29. As highlighted in Sun et al. (2023, Section 2), the modified half-normal distribution appears in several Bayesian statistical methods as a posterior distribution to sample from in Bayesian Binary regression, analysis of directional data, and Bayesian graphical models.
@@
which shows that Assumption 2-(ii) is verified for any $K\geq 0$, and Assumption 2-(iii) is verified for $\mu=2\xi$. Therefore, we can conclude that 29 satisfies Assumption 2.
@@ 3.4.2 Multidimensional Gaussian mixture distribution
We consider a multidimensional Gaussian mixture data distribution with unknown mean and variance, i.e.,
$\pi_{\mathsf{D}}(\mathrm{d}x)=\sum_{i=1}^{I}\widetilde{\xi}_{i}\frac{1}{(2\pi s_{i}^{2})^{d/2}}\exp\left(-\frac{|x-\eta_{i}|^{2}}{2s_{i}^{2}}\right)\ \mathrm{d}x,\quad x\in\mathbb{R}^{d},$ (31)
with $s_{i}>0$, $\eta_{i}\in\mathbb{R}^{d}$, and $\widetilde{\xi}_{i}\in[0,1]$ for $i\in\{1,\dots,I\}$ such that $\sum_{i=1}^{J}\widetilde{\xi}_{i}=1$. The authors in Silveri & Ocello (2025, Appendix A) show that the score function of 31 is Lipschitz continuous and $\log\pi_{\mathsf{D}}$ is weakly convex. Therefore, Assumption 2 is satisfied. In addition, the distribution 31 covers also case of the double-well potential:
$U(x)=|x|^{4}-|x|^{2},\quad x\in\mathbb{R}^{d},$ (32)
@@ 3.4.3 Multi-dimensional Potentials
verifies Assumption 2 with $K=0$, $R=1$, and $\mu=2$ as well as the following non-convex potential presented in Johnston et al. (2025, Example 4.2):
@@ 4 Related Work and Comparison
The first approach focuses on $\alpha$-divergences, particularly the Kullback–Leibler (KL) divergence and Total Variation (TV) distance (e.g., Benton et al. (2024); Conforti et al. (2025); Yang & Wibisono (2023); Li & Cai (2024); Block et al. (2020); De Bortoli et al. (2021); Lee et al. (2022); Li et al. (2024); Lee et al. (2023); Chen et al. (2023a; b); Oko et al. (2023); Liang et al. (2025); Yang & Wibisono (2022)), which are the vast majority of the results available in the literature. Crucially, bounds on KL divergence imply bounds on TV distance via Pinsker’s inequality, strengthening their wide applicability. We provide a brief and selective overview of some of the findings following this first approach. The results in TV distance in Lee et al. (2022) and in KL divergence Yang & Wibisono (2023) established convergence bounds characterized by polynomial complexity under the assumption that the data distribution satisfies a logarithmic Sobolev inequality and that the score function is Lipschitz continuous. By replacing the requirement that the data distribution satisfies a functional inequality with the assumption that $\pi_{\text{D}}$ has finite KL divergence with respect to the standard Gaussian and by assuming that the score function for the forward process is Lipschitz, the authors in Chen et al. (2023b) managed to derive bounds in TV distance which scale polynomially in all the problem parameters. By requiring only the Lipschitzness of the score at the initial time rather than along the full trajectory, the authors in Chen et al. (2023a, Theorem 2.5) managed to establish, using an exponentially decreasing then linear step size, convergence bounds in KL divergence with quadratic dimensional dependence and logarithmic complexity in the Lipschitz constant. Later, Benton et al. (2024) provided KL convergence bounds that are linear in the data dimension, up to logarithmic factors, by assuming finite second moments of the data distribution and employing early stopping. However, both the results of Chen et al. (2023a, Theorem 2.5) and Benton et al. (2024, Theorem 1 and Corollary 1) still require the uniqueness of solutions for the backward SDE 2, and therefore additional assumptions on the score function are needed. For further discussion on this point, we refer the reader to Bruno et al. (2025, Section 4.2). Assuming finite second moments and using an exponential integrator (EI) scheme with both constant and exponentially decaying step sizes, the authors in Conforti et al. (2025, Corollary 2.4) derive a KL divergence bound with early stopping, which scales linearly in the data dimension up to logarithmic factors. Bounds in KL without early stopping have been derived in Conforti et al. (2025) for data distributions with finite Fisher information with respect to the standard Gaussian distribution. We note that this condition on $\pi_{\text{D}}$ stated in Conforti et al. (2025, Assumption H2) still requires that the potential $U\in C^{1}(\mathbb{R}^{d})$. The KL bounds provided in Conforti et al. (2025, Theorem 2.1 and 2.2) scale linearly in the Fisher information when an EI discretization scheme with constant step size is used, and logarithmically in the Fisher information when an exponential-then-constant step size Conforti et al. (2025, Theorem 2.3) is employed.
@@
The second approach focuses on convergence bounds in Wasserstein distance, a metric which is often considered more practical and informative for estimation tasks. We can relate results following this approach with the results of the first approach only when $\pi_{\mathsf{D}}$ is a strongly log-concave distribution. In this case, $W_{2}$-bounds in terms of KL divergence follow from an extension of Talagrand’s inequality (Gozlan & Léonard, 2010, Corollary 7.2). However, for two general data distributions, there is no known relationship between their KL divergence and their $W_{2}$. Therefore, we cannot compare our findings in Theorem 19 and Theorem 21 with the results derived following the first approach. One line of work within the second approach assumes (at least) strong log-concavity of the data distribution (Strasman et al., 2025; Gao et al., 2025; Bruno et al., 2025; Tang & Zhao, 2024; Yu & Yu, 2025). Under this (strict) assumption, Bruno et al. (2025, Remark 12) achieved optimal data dimensional dependence, i.e., reaching $O(\sqrt{d})$. The recent bound in Silveri & Ocello (2025, Theorem D.1) exhibits similar scaling in $d$ while relaxing the strong log-concavity assumption on $\pi_{\mathsf{D}}$ to weakly log-concavity, but still requiring that the potential $\nabla^{2}U$ exists (see, e.g., Silveri & Ocello (2025, Proof of Proposition B.1 and B.2)). Our Assumption 2 is much weaker than this requirement and it allows to consider the case of potentials with discontinuous gradients covering a wider range of distributions as outlined in Section 3.4. Another line of work following this approach focuses on specific structural assumptions of the data distribution. For instance, convergence bounds in Wasserstein distance of order one with exponential dependence on the problem parameters have been obtained in Bortoli (2022) under the so-called manifold hypothesis, namely assuming that the target distribution is supported on a lower-dimensional manifold or is given by some empirical distribution. Under the same metric, the authors in Mimikos-Stamatopoulos et al. (2024) provide a convergence analysis when the data distribution is defined on a torus. Under the $W_{2}$ metric, Wang & Wang (2024) derive convergence bounds assuming that the tail of $\pi_{\mathsf{D}}$ is Gaussian and that $U\in C^{2}$, which is a stronger condition than merely requiring $\nabla U$ to be Lipschitz. We summarize in Table 1 and the best results obtained in $W_{2}$, i.e., Bruno et al. (2025); Silveri & Ocello (2025) and compare with our best result, which scale polynomially in the data dimension, i.e. $O(\sqrt{d})$ in Theorem 19. As mentioned in the Introduction, previous $W_{2}$ bounds (Silveri & Ocello, 2025; Gao et al., 2025; Strasman et al., 2025; Tang & Zhao, 2024) require the stepsize $\gamma$ of the generative algorithm to be tuned based on quantities that are often difficult to compute in practice, such as the Lipschitz or strong convexity constant of the data distribution, which can lead to very small stepsizes. In contrast, Theorem 19 and Theorem 21 impose no such restrictions, making them more suitable for practical use. Table 2 summarizes the assumptions on $\gamma$ in prior works and compares them with ours.
We close this section by briefly commenting on the choice of deriving our results in Wasserstein distance of order two. Beyond its theoretical relevance, this choice is motivated by practical considerations in generative modeling. First, the Wasserstein distance is often regarded as a more informative and robust metric for estimation tasks. Second, a widely used performance metric for evaluating the quality of images produced by generative models is the Fréchet Inception Distance (FID) Heusel et al. (2017), which measures the Fréchet distance between the distributions of generated and real samples, assuming Gaussian distributions. In particular, this Fréchet distance is equivalent to the Wasserstein-2 distance. Thus, providing convergence results under the Wasserstein-2 metric enhances the practical relevance of our theoretical findings.
@@
Table 1: Summary of previous bounds for $W_{2}(\mathcal{L}(\widehat{Y}_{J}^{\text{EM}}),\pi_{\mathsf{D}})$ and our result in Theorem 19. All the bounds assume that $\pi_{\text{D}}(\mathrm{d}x)\propto e^{-U(x)}\mathrm{d}x$ has finite second moments.
@@
$U$ strongly convex, $\nabla\log p_{t}(0)\in L^{2}([\epsilon,T])$, and Assumption 4 $O(\sqrt{d})\sqrt{\epsilon}+O(\sqrt{d})e^{-2\widehat{L}_{\text{MO}}(T-\epsilon)-\epsilon}+O(e^{(1+\zeta-2\widehat{L}_{\text{MO}})(T-\epsilon)})\sqrt{\varepsilon_{\text{SN}}}+O(\sqrt{d}e^{T^{2\alpha+1}}T^{2\alpha+1}\widetilde{\varepsilon}^{1/2}_{\text{AL}})\gamma^{1/2},$ with $\widehat{L}_{\text{MO}}>0$ lower bound of the strongly convex constant of $U$, see e.g., Bruno et al. (2025, Remark 4). Bruno et al. (2025, Remark 12)
$U\in C^{2}(\mathbb{R}^{d})$, weakly convex, and Assumption 4 $e^{(2L_{U}+5)\eta(\beta,L,(2L_{U}+5)^{2}\gamma/2)}[e^{-T}W_{2}(\pi_{\mathsf{D}},\pi_{\infty})+4\varepsilon_{\text{SN}}(T-\eta(\beta,L,0))+\sqrt{2\gamma}(4L_{U}\sqrt{d}+6\sqrt{d}+\sqrt{d+\mathbb{E}[|X_{0}|^{2}]})(T-\eta(\beta,L,0))]$, with $L_{U}\geq 0$ one-sided Lipschitz constant for $\nabla U$, see e.g., Silveri & Ocello (2025, Assumption H1), $\eta(\beta,L,\gamma)$ defined in (Silveri & Ocello, 2025, equation (29)), and $\gamma<2/(2L_{U}+5)^{2}$ . Silveri & Ocello (2025, Theorem D.1)
Assumption 2 and Assumption 4 O(d) ϵ + O(d) e^ - 2 ∫_ϵ^T β^OS, K, μ_t dt -ϵ + O(e^(1+ ζ)(T-ϵ) - 2 ∫_ϵ^T β^OS, K, μ_t dt ) ε_SN + O(d e^T^2α+1 T^3 α+1 ~ε^1/2_AL) γ^1/2. Theorem 19
Table 2: Summary of restrictions on the stepsizes of the generative algorithm $\widehat{Y}_{J}^{\text{EM}}$ used in the previous bounds for $W_{2}(\mathcal{L}(\widehat{Y}_{J}^{\text{EM}}),\pi_{\mathsf{D}})$ and our results in Theorem 19 and Theorem 21. All the bounds assume that $\pi_{\text{D}}(\mathrm{d}x)\propto e^{-U(x)}\mathrm{d}x$ has finite second moments.
Assumption on $\pi_{\text{D}}$ Restriction on the stepsize Reference
$U$ strongly convex, $\nabla U$ Lipschitz continuous, and Assumption 4 0< γ≤min_0 ≤t ≤T ( 1-e-2t(1m0-1)(1+e-2t(1m0-1))(1+4 (~L(t))2+ 2M1) ) and 0< γ≤min_0 ≤t ≤T ( 1+(1m0-1)e-2t1 - (1m0-1)e-2t), with $m_{0}>0$ strong convexity constant for $U$, see e.g., Gao et al. (2025, Assumption 1), $M_{1}>0$ defined in Gao et al. (2025, Assumption 2), $\widetilde{L}(t)=\min_{0\leq t\leq T}((1-e^{-2t})^{-1},e^{2t}L_{0})$ Lipschitz constant of $\nabla\log p_{t}(x)$, and $L_{0}>0$ Lipschitz constant for $\nabla U$, see e.g., Gao et al. (2025, Assumption 1). Gao et al. (2025, Assumption 4) used in Gao et al. (2025, Theorem 2)
$U$ strongly convex, $\nabla\log p_{t}(0)\in L^{2}([\epsilon,T])$, and Assumption 4 γ∈(0,1). Bruno et al. (2025, Remark 12)
$U$ strongly convex, and $\nabla U$ Lipschitz continuous 0< γ< C(T-t)( maxtj≤s ≤tj+1L(T-s) ) L(T-t) e^-(t_j+1-t_j), with $\left\{t_{j},\ 0\leq j\leq J\right\}$ regular discretization of $[0,T]$, $C_{T-t}$ and $L_{T-t}$ are strong log-concavity and Lipschitz constant, respectively for $\nabla\log\frac{p_{t}}{\varphi_{\sigma^{2}}}$, where $\varphi_{\sigma^{2}}$ is the density function of a mean zero Gaussian distribution with variance $\sigma^{2}I_{d}$. Strasman et al. (2025, Proposition C.3) used in Strasman et al. (2025, Theorem 4.2)
$U$ strongly convex, $\nabla U$ Lipschitz continuous, and Assumption 4 0< γ< min( 12, κ2 T(1+ κ) ), with $\kappa>0$ strong convexity constant. Tang & Zhao (2024, Theorem 3)
$U\in C^{2}(\mathbb{R}^{d})$, weakly convex, and Assumption 4 0< γ< 2(2LU+5)2, with $L_{U}\geq 0$ one-sided Lipschitz constant for $\nabla U$, see e.g., Silveri & Ocello (2025, Assumption H1). Silveri & Ocello (2025, Theorem D.1)
Assumption 2, and Assumption 4 γ∈(0,1). Theorem 19, and Theorem 21.